40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1298     Entry time: Thu Feb 12 17:43:33 2009
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: SRC strangeness solved 
I found the problem with the DRMI lock I had last night was caused by the zero gain in the PD11_I filter.
I don't know how it happened but putting it back to 1.000 made the DRMI lock far more stable and AS166Q got more than 3000.

I also re-centered POY PD to remove the offset in the y-arm loop. The large power drops while y-arm is locked by itself were eliminated.
ELOG V3.1.3-